Solve for the exact solutions in the interval [0,2π)
cos(2x)=cos(x)
I got 0 but apparently there are more solutions


Answer by Alan3354(69443)

cos(2x)=cos(x)
-----
2cos^2(x) - 1 = cos(x)
2cos^2(x) - cos(x) - 1 = 0
(2cos + 1)*(cos - 1) = 0
cos = 1
x = 0
-----
cos = -1/2
x = 2pi/3
x = 4pi/3